# Enhancing the ODMR Signal of Organic Molecular Qubits

### v2

@@ Abstract
In the fields of quantum information science and sensing, electron spins are often purified into a specific polarisation through an optical-spin interface, a process known as optically-detected magnetic resonance (ODMR). Diamond-NV centres and transition metals are both excellent platforms for these so-called colour centres, while metal-free molecular analogues are also gaining popularity for their extended polarisation lifetimes, milder environmental impacts, and reduced costs. In our earlier attempt at designing such organic high-spin $\pi$-diradicals, we proposed to spin-polarise by shelving triplet $M_{S}=\pm 1$ populations as singlets. This was recently verified by experiments albeit with low ODMR contrasts of $<1\%$ at temperatures above 5 K. In this work, we propose to improve the ODMR signal by moving singlet populations back into the triplet $M_{S}=0$ sublevel, designing a true carbon-based molecular analogue to the NV centre. To achieve both spin channels in $\pi$-diradicals, we leverage on weaker spin-orbit couplings beyond the nearest-neighbour approximation, made possible by careful control of orbital nodal structures and group-theoretical considerations. These analyses are further confirmed by ab initio calculations of a realistic trityl-based radical dimer. Microkinetic analyses point towards high ODMR contrasts of around $30\%$ under experimentally-feasible conditions, a stark improvement from previous works. Finally, in our quest towards ground-state optically-addressable molecular spin qubits, we exemplify how our symmetry-based design avoids Zeeman-induced singlet-triplet mixings during electron paramagnetic resonance (EPR) experiments, setting the scene for realising electron spin qubit gates.
@@ Introduction
As a first step along this path, organic molecules were designed to demonstrate ODMR in high-spin excited states while acknowledging the limited lifetimes of these electronic excitations [43, 44, 45, 46, 47, 48, 49, 50]. Subsequently, our group theoretically proposed a class of organic $\pi$-diradicals with optically-addressable high-spin ground states [51] that were also experimentally validated by Chowdhury et al. [52] and Kopp et al. [53]. In all of these systems, spin polarisation was likely attained by shelving the triplet $M_{S}=\pm 1$ population in the singlet manifold via spin-selective intersystem crossings (ISCs) while utilising the presence of additional singlet charge-recombined states absent from the triplet manifold [Fig. Introductionb]. (Interestingly, such imbalances have recently been proposed as evidence of low-energy electronic spin isomers [54].) Missing from the ODMR cycles are the ISCs between singlets and triplets selective for the $M_{S}=0$ triplet sublevel, which the NV centre exhibits. As mentioned in our previous work [51], this limits the maximum possible triplet $M_{S}=0$ population to $25\%$, constraining the ODMR resolution. Moreover, this increases the risk of metastable singlet molecules returning as $M_{S}=\pm 1$ triplets, which would negate the accumulated spin polarisation. Indeed, in the aforementioned experimental works, poor ODMR contrasts of $<1\%$ were observed at temperatures above 5 K. We note in passing that, apart from $\pi$-diradicals, alkaline earth metal complexes [55, 56] and nitrenes [57] also display diradical properties and their functions as molecular colour centres have not been fully explored.
@@
Here, we design another class of organic $\pi$-diradicals, this time capable of ISCs involving both triplet $M_{S}=\pm 1$ and $M_{S}=0$ sublevels of different electronic states [Fig. Introductionc]. This way, we create a true analogue of the NV centre, which we predict to demonstrate enhanced ODMR signals. We are motivated by the presence of beyond-nearest-neighbour (through-space) SOCs that were not explored in our previous nearest-neighbour-only (through-bond-only) semi-empirical model [51]. As explained in the Supplementary Information of that work, such SOCs can facilitate ISCs among the lowest triplet local excitations (LEs) and the lowest singlet charge transfers (CTs) [Fig. Introductionb], whereby the LEs comprise linear combinations of each monoradical’s lowest electronic transitions (hence “local”) while the CTs represent charge-recombination states of the two radicals (also known as zwitterionic states). This effect was subsequently observed by Kopp et al. [53] (albeit in the absence of a $\Delta M_{S}=0$-selective ISC that limited the ODMR contrast), simplifying the electronic structure analysis to only the first four singlet and first two triplet excitations [Fig. Introductionb]. Preliminary analysis conducted by the present work found only modest spin selectivities of $68.1\%$ from next-nearest-neighbour SOCs, in contrast to the $100\%$ selectivity from nearest-neighbour SOCs endowed by symmetry. Subsequent optimisation leveraging on interferences from higher-nearest-neighbour SOCs enhances the former value to $100\%$ with opposite spin selectivites from their nearest-neighbour counterparts. Our design principle stems from a careful group-theoretical analysis of the transitioning orbitals, which coincidentally also solves the problem of singlet-triplet mixing during electron paramagnetic resonance (EPR) measurements of weakly-coupled diradicals (as most $\pi$-diradicals are [58, 53]). As a realistic prototypical example of our design, the electronic structure of PT2TM-p-PT2TM [Fig. Introductionc] was shown at both density functional theory (DFT) and multi-configurational (MC) levels of theory to demonstrate an ODMR mechanism similar to the NV centre. The potential for signal enhancement through this ODMR pathway is supported by microkinetic analyses, where we found high ODMR contrasts of around $30\%$ to be experimentally possible (as compared to $<1\%$ in earlier studies [52, 53]) While not explored by the present work, we expect further mesitylation [Fig. Introductionc] to improve the luminescence quantum yield via geometrical relaxation [59, 60], as exemplified by Chowdhury et al. [52] to produce near-unity photoluminescence yields. This paves the way towards better designs of optically-addressable organic spins. Note that the alternacy symmetry of PT2TM-p-PT2TM is not important to our design, which should be separated from our earlier work [51].
@@
FIG. 1: : (a) Simplified ODMR mechanism typical of diamond-NV defects. (b) Simplified ODMR mechanism of alternant $\pi$-diradicals as explored by our previous work [51]. Note that the CT-to-GS ISC rates are small if the two benzylic radicals are tethered at their meta positions, which is avoided by our new design (see main text). (c) ODMR mechanism of our new design, which more closely resembles the spin polarisation pathway of NV centres. Also drawn are the PT2TM-p-PT2TM diradical and its mesitylated counterpart, both of which satisfy our design requirements. (Abbreviations of electronic transitions: Abs = absorption; PL = photoluminescence; IC = internal conversion; ISC = intersystem crossing. Labels for electronic states: GS = ground state; LE = local excitation; CT = charge transfer. These labels are inherited from our previous analysis of the $\pi$-diradical electronic structure [51] with subscripts signifying the state irreps.)
@@
As we have previously shown [51], when two benzylic radicals are covalently connected with a significant torsion (as is the way most $\pi$-diradicals are experimentally constructed [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 58]), its SOC matrix elements are dominated by nearest-neighbour interactions between the two $2p$ atomic orbitals (AOs) of opposite $\pi$-rings (reason: El-Sayed rules) that are closest to the dimer linkage (reason: SOC is a local effect). Henceforth, we refer solely to $\pi$ AOs. Furthermore, if the spin quantisation axis lies parallel (perpendicular) to the line connecting the two $2p$ AO centres, then the ISC spin selectivity arising from the SOC matrix elements is strictly $\Delta M_{S}=0$ (±$\pm 1$) due to symmetry [51]. Note that the last condition is satisfied for interactions at the dimer bond because the spin axes (wihout applying any magnetic fields) coincide with the dimer’s high symmetry axes. However, ISCs in $\pi$-diradicals occur not between pairs of $2p$ AOs but between pairs of delocalised molecular orbitals (MOs) that are each a linear combination of a few $2p$ AOs. When one of the transitioning MOs contains a node at the dimer linkage, the relevant SOC matrix elements necessarily involve next-nearest-neighbour couplings not considered by the above analysis. Such effects are generally small following the local nature of SOCs, yet they can be appreciable if most of the transitioning MO’s amplitude is piled up at these next-nearest-neighbour sites (as opposed to nearest-neighbour couplings weighted by small MO amplitudes). Importantly, next-nearest-neighbour interactions involve two $2p$ AO centres that are not colinear with the spin quantisation axis, causing the ISC selectivity to break down [Fig. Investigating the role of beyond-nearest-neighbour SOCsa]. Indeed, for two benzylic radicals covalently tethered at the para positions with $90^{\text{o}}$ torsion (hereafter called benzyl-p-benzyl), the next-nearest-neighbour SOC evaluated over singly-occupied $2p$ AOs biases the $\Delta M_{S}=\pm 1$ transitions with $68.1\%$ selectivity (noting that the spin quantisation ($z$-)axis lies parallel to the dimer linkage due to the molecular $D_{2d}$ point group symmetry) and this SOC has a significant amplitude of $0.13\mathrm{\,cm^{-1}}$ [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. For comparsion, the nearest-neighbour analogue has a perfect spin selectivity of $\Delta M_{S}=0$ with a larger amplitude of $1.37\mathrm{\,cm^{-1}}$. Note that these SOC matrix elements represent upper bounds to the true molecular values because we have assumed one full electron in each AO.
@@
That ISC in $\pi$-diradicals can actually occur with both spin selectivities is contrary to most theoretical analyses [78, 79, 80] (including ours [51]) and opens up the possibility of transferring the full electronic structure of NV centres into molecules. Our next step is to perfect the next-nearest-neighbour spin selectivity via symmetry considerations. We note that by connecting two benzylic radicals at their para positions, the resulting diradical has at least a $D_{2}$ point group symmetry with irreducible representations (irreps) organised by their characters under $C_{2}\left(y\right)$ and $C_{2}\left(z\right)$ rotations. The irrep of any electronic state is in turn the direct product of its orbital and spin irreps. Focusing on the orbital part, if we continue to align the $z$-axis along the dimer linkage [Fig. Investigating the role of beyond-nearest-neighbour SOCsb,e], then a $C_{2}\left(y\right)$ rotation becomes equivalent to swapping the two monomers. Therefore, for two weakly-interacting $\pi$-systems of significant torsion, the low-lying excitations are expected to appear in near-degenerate pairs of $\left\{A,B_{1}\right\}$ and $\left\{B_{2},B_{3}\right\}$, each comprising a state of opposite parity under $C_{2}\left(y\right)$, i.e. equal probability of exciting either monomer. These two pairs can be distinguished by their transformations under $C_{2}\left(z\right)$ and, for that, it is easier to think in the basis of monomeric $\pi$ MOs. We first note that $2p$ AOs have $\pi$ symmetry, that is, their phases change under $C_{2}$ rotation about the AO centre. Further, for the two $2p$ AOs at the dimer linkage, that $\pi$-symmetry-defining $C_{2}$ rotation coincides with the diradical’s $C_{2}\left(z\right)$ symmetry element. Hence, a monomeric MO with a non-zero amplitude at the dimer linkage must be antisymmetric under $C_{2}\left(z\right)$ rotation. By contrast, for a monomeric MO to be symmetric under $C_{2}\left(z\right)$ rotation, it must possess a node at the dimer linkage. For instance, the Hückel method predicts the highest occupied, singly occupied and lowest unoccupied MOs (abbreviated as HOMO, SOMO, and LUMO respectively) of the benzylic radical to have $+1$, $-1$, and $+1$ characters under $C_{2}\left(z\right)$ rotation [Fig. Investigating the role of beyond-nearest-neighbour SOCsc]. Finally, we recall that the symmetry of the orbital wavefunction is the direct product of the symmetries of all singly occupied orbitals. Therefore, the ground states of benzyl-p-benzyl, having two unpaired electrons in the SOMOs, transform as either $A$ or $B_{1}$. Similarly, the lowest CT singlets with all doubly-occupied orbitals also transform as either $A$ or $B_{1}$. As for the lowest LEs, they are a linear combination of HOMO-to-SOMO and SOMO-to-LUMO excitations (an outcome of alternancy symmetry; see Ref. [51]) and hence transform as either $B_{2}$ or $B_{3}$. Including the characters under $C_{2}\left(y\right)$ rotation yields the approximate energy level diagram illustrated in Fig. Investigating the role of beyond-nearest-neighbour SOCsd (for a review of the electronic structure of benzylic diradicals, see our earlier work [51]).
@@
Returning to our goal of attaining perfect ISC spin selectivities, we note that the singlet, triplet $M_{S}=0$, and triplet $M_{S}=\pm 1$ spin sublevels transform as $A$, $B_{1}$, and $\left\{B_{2},B_{3}\right\}$ respectively [81, 82]. For an ISC to be symmetry-allowed, the SOC matrix elements must contain the totally symmetric irrep. Hence, because the SOC Hamiltonian is totally symmetric, ISC processes in benzyl-p-benzyl are spin-selective for $\Delta M_{S}=\pm 1$ when moving between triplet LEs and singlet CTs, and $\Delta M_{S}=0$ when moving between singlet CTs and triplet GSs. Actually, we already knew this from the previous section: The first ISC occurs between SOMOs and LUMOs (or HOMOs) of opposite monomers and only the latter has a node at the dimer linkage [Fig. Investigating the role of beyond-nearest-neighbour SOCsc]. Thus, this process occurs predominantly via next-nearest-neighbour SOCs, which we found before to exhibit $\Delta M_{S}=\pm 1$ selectivity. Similarly, the second ISC connects between two SOMOs of opposite monomers and is mediated mostly by nearest-neighbour SOCs with $\Delta M_{S}=0$ selectivity. Yet, the key improvement achieved by this section is the perfection of ISC spin selectivities to $100\%$ by making the opposite spin channel symmetry-forbidden. From a microscopic perspective, we have destructively interfered any deleterious ISC channels using other higher-nearest-neighbour SOCs, treated more generally by the group-theoretical approach. Importantly, this electronic structure produces the ODMR mechanism shown in Fig. Introductionc, where populations in the triplet $M_{S}=\pm 1$ LEs are moved into the triplet $M_{S}=0$ GS via two sequential ISCs of opposite spin selectivities. In other words, such molecules exhibit the ODMR mechanism of NV centres. As an aside, while it is tempting to consider only symmetry arguments and ignore the discussions in the previous section, note that symmetry-allowed transitions need not occur with high amplitudes. Therefore, one still needs to check if the allowed ISCs are facilitated by either nearest- or next-nearest-neighbour SOCs, which have the largest amplitudes. For example, the meta-connected diradicals explored in our previous work [51] have negligible SOCs between singlet CTs and triplet GSs [Fig. Introductionb] because they are dominated by third-nearest-neighbour effects with a maximum SOC amplitude of $0.02\mathrm{\,cm^{-1}}$ [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. In other words, our plan is made viable by our specially-chosen para-connectivity – a consequence of our group-theoretical analysis (meta-connectivity does not yield $D_{2}$ point group symmetry).
@@
As speculated, the lowest LE transitions are largely localised on the central (tetra-chlorinated) benzyl-p-benzyl fragment [Fig. Our proposalc]. Hence, for computational efficiency, the SOC matrix elements were computed for this central fragment instead, capping any open ends with hydrogen. Indeed, the MCSCF/CI results suggest appreciable SOC matrix elements along the proposed ODMR pathway [Fig. Our proposalb]. While, strictly speaking, these calculations present only an upper bound to the true SOC matrix elements of PT2TM-p-PT2TM, their values should not decrease by much upon weak $\pi$-conjugation to the additional trichlorophenyl substituents. Most importantly, these calculations lend credence to the theoretical proposal presented by this work.
@@
The alternative approach would be to somehow make elements that differ between ${\bf g}_{1}$ and ${\bf g}_{2}$ disappear under experimental conditions. This can be achieved in our diradicals by aligning the magnetic field ${\bf B}$ parallelly to the dimer linkage. To see that, we first note that elements of the $g$-tensor have the following approximate expression, obtained from second-order perturbation theory in the SOC: [88, 89]
@@
The singlet-triplet states may now block-diagonalise $\hat{H}_{\text{spin}}$ with no singlet-triplet mixings present. Furthermore, due to the approximate $D_{2d}$ point group symmetry of our system, the zero-field splitting caused by ${\bf D}$ also occurs along the $z$-axis. This may be further rationalised by the cylindrical spin density distribution of the diradical [89]. Therefore, the singlet-triplet basis fully-diagonalises $\hat{H}_{\text{spin}}$ in Eq. (23).
@@
The kinetic processes connecting these eigenstates and their associated rate constants are presented in Fig. Predicted ODMR contrastsa; there, the symbols adopt their usual meanings. As far as possible, model parameters were selected within experimental reason. Of particular importance are the following:
@@
FIG. 4: : (a) Schematic diagram depicting the excited-state transitions being considered by our microkinetic model. (b) Steady-state solutions to the optically-induced spin polarisation, ODMR contrast, triplet $M_{S}=0$ GS population, and triplet $M_{S}=0$ LE population (which simulates the emission intensity). Plots are made across varying optical pump rates $k_{\text{abs}}$ and spin decoherence rates $k_{\text{decoh}}$ within experimental ranges. Other parameters are defined in (a), chosen to be as realistic as possible.

### v3

@@ Abstract
In quantum information science and sensing, electron spins are often purified into a specific polarisation through an optical-spin interface, a process known as optically-detected magnetic resonance (ODMR). Diamond-NV centres and transition metals are both excellent platforms for these so-called colour centres, while metal-free molecular analogues are also gaining popularity for their extended polarisation lifetimes, milder environmental impacts, and reduced costs. In our earlier attempt at designing such organic high-spin $\pi$-diradicals, we proposed to spin-polarise by shelving triplet $M_{S}=\pm 1$ populations as singlets. This was recently verified by experiments albeit with low ODMR contrasts of $<1\%$ at temperatures above 5 K. In this work, we propose to improve the ODMR signal by moving singlet populations back into the triplet $M_{S}=0$ sublevel, designing a true carbon-based molecular analogue to the NV centre. Our proposal is based upon transition-orbital and group-theoretical analyses of beyond-nearest-neighbour spin-orbit couplings, which are further confirmed by ab initio calculations of a realistic trityl-based radical dimer. Microkinetic analyses point towards high ODMR contrasts of around $30\%$ under experimentally-feasible conditions, a stark improvement from previous works. Finally, in our quest towards ground-state optically-addressable molecular spin qubits, we exemplify how our symmetry-based design avoids Zeeman-induced singlet-triplet mixings, setting the scene for realising electron spin qubit gates.
@@ Introduction
As a first step along this path, organic molecules were designed to demonstrate ODMR in high-spin excited states while acknowledging the limited lifetimes of these electronic excitations [43, 44, 45, 46, 47, 48, 49, 50, 51]. Subsequently, our group theoretically proposed a class of organic $\pi$-diradicals with optically-addressable high-spin ground states [52] that were also experimentally validated by Chowdhury et al. [53] and Kopp et al. [54]. In all of these systems, spin polarisation was likely attained by shelving the triplet $M_{S}=\pm 1$ population in the singlet manifold via spin-selective intersystem crossings (ISCs) while utilising the presence of additional singlet charge-recombined states absent from the triplet manifold [Fig. Introductionb]. (Interestingly, such imbalances have recently been proposed as evidence of low-energy electronic spin isomers [55].) Missing from the ODMR cycles are the ISCs between singlets and triplets selective for the $M_{S}=0$ triplet sublevel, which the NV centre exhibits. As mentioned in our previous work [52], this limits the maximum possible triplet $M_{S}=0$ population to $25\%$ at steady state, constraining the ODMR resolution. Moreover, this increases the risk of metastable singlet molecules returning as $M_{S}=\pm 1$ triplets, which would negate the accumulated spin polarisation. Indeed, in the aforementioned experimental works, poor ODMR contrasts of $<1\%$ were observed at temperatures above 5 K. We note in passing that, apart from $\pi$-diradicals, alkaline earth metal complexes [56, 57] and nitrenes [58] also display diradical properties and their functions as molecular colour centres have not been fully explored.
@@
Our earlier theoretical model [52] had focused primarily on nearest-neighbour-only (through-bond-only) SOCs, which predicted ISCs to have $\Delta M_{S}=0$ (±$\pm 1$) selectivity for electron spins quantised parallelly (perpendicularly) to the dimer linkage. However, beyond-nearest-neighbour SOC effects have been shown to be experimentally relevant in ODMR, as in the recent work by Kopp et al. [54] [Fig. Introductionb], and such effects can induce ISCs with an opposite spin selectivity. Motivated by this observation, in this work, we design another class of organic $\pi$-diradicals, this time capable of ISCs involving both triplet $M_{S}=\pm 1$ and $M_{S}=0$ sublevels of different electronic states as facilitated by beyond-nearest-neighbour and nearest-neighbour SOCs respectively [Fig. Introductionc]. This way, we create a true analogue of the NV centre, which we predict to demonstrate enhanced ODMR signals by optically pumping the near-degenerate triplet and singlet local excitations. Our design principle stems from a careful group-theoretical analysis of the transitioning orbitals, which coincidentally also solves the problem of singlet-triplet mixing during electron paramagnetic resonance (EPR) measurements of weakly-coupled diradicals (as most $\pi$-diradicals are [59, 54]). As a realistic prototypical example of our design, the electronic structure of PT2TM-p-PT2TM [Fig. Introductionc] was shown at both density functional theory (DFT) and multi-configurational (MC) levels of theory to demonstrate an ODMR mechanism similar to the NV centre. The potential for signal enhancement through this ODMR pathway is supported by microkinetic analyses, where we found high ODMR contrasts of around $30\%$ to be experimentally possible (as compared to $<1\%$ in earlier theoretical [52] and experimental [53, 54] studies). In this regime, the steady-state triplet population exceeds $25\%$, which is the upper bound of prior works [52, 53, 54]. While not explored by the present work, we expect further mesitylation [Fig. Introductionc] to improve the luminescence quantum yield via geometrical relaxation [60, 61], as exemplified by Chowdhury et al. [53] to produce near-unity photoluminescence yields. This paves the way towards better designs of optically-addressable organic spins. Note that the alternacy symmetry of PT2TM-p-PT2TM is not important to our design, which should be separated from our earlier work [52].
@@
FIG. 1: : (a) Simplified ODMR mechanism typical of diamond-NV defects. (b) Simplified ODMR mechanism of alternant $\pi$-diradicals as explored by our previous work [52]. Note that the CT-to-GS ISC rates are small if the two benzylic radicals are tethered at their meta positions, which is avoided by our new design (see main text). (c) ODMR mechanism of our new design, which more closely resembles the spin polarisation pathway of NV centres. Also drawn are the PT2TM-p-PT2TM diradical and its mesitylated counterpart, both of which satisfy our design requirements. (Abbreviations of electronic transitions: Abs = absorption; PL = photoluminescence; IC = internal conversion; ISC = intersystem crossing. Labels for electronic states: GS = ground state; LE = local excitation; CT = charge transfer. These labels are inherited from our previous analysis of the $\pi$-diradical electronic structure [52] with subscripts $A$, $B_{1}$, $B_{2}$, and $B_{3}$ signifying the state irreps.)
@@
As we have previously shown [52], when two benzylic radicals are covalently connected with a significant torsion (as is the way most $\pi$-diradicals are experimentally constructed [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 59]), its SOC matrix elements are dominated by nearest-neighbour interactions between the two $2p$ atomic orbitals (AOs) of opposite $\pi$-rings (reason: El-Sayed rule  [^1] The El-Sayed rule states that the SOC matrix elements are the largest when the transitioning orbitals adopt different orientations [El-Sayed1963].) that are closest to the dimer linkage (reason: SOC is a local effect). Henceforth, we refer solely to $\pi$ AOs. Furthermore, if the spin quantisation axis lies parallel (perpendicular) to the line connecting the two $2p$ AO centres, then the ISC spin selectivity arising from the SOC matrix elements is strictly $\Delta M_{S}=0$ (±$\pm 1$) due to symmetry [52]. Note that the last condition is satisfied for interactions at the dimer bond because the spin axes (wihout applying any magnetic fields) coincide with the dimer’s high symmetry axes. However, ISCs in $\pi$-diradicals occur not between pairs of $2p$ AOs but between pairs of delocalised molecular orbitals (MOs) that are each a linear combination of a few $2p$ AOs. When one of the transitioning MOs contains a node at the dimer linkage, the relevant SOC matrix elements necessarily involve next-nearest-neighbour couplings not considered by the above analysis. Such effects are generally small following the local nature of SOCs, yet they can be appreciable if most of the transitioning MO’s amplitude is piled up at these next-nearest-neighbour sites (as opposed to nearest-neighbour couplings weighted by small MO amplitudes). Importantly, next-nearest-neighbour interactions involve two $2p$ AO centres that are not colinear with the spin quantisation axis, causing the ISC selectivity to break down [Fig. Investigating the role of beyond-nearest-neighbour SOCsa]. Indeed, for two benzylic radicals covalently tethered at the para positions with $90^{\text{o}}$ torsion (hereafter called benzyl-p-benzyl), the next-nearest-neighbour SOC evaluated over singly-occupied $2p$ AOs biases the $\Delta M_{S}=\pm 1$ transitions with $68.1\%$ selectivity (noting that the spin quantisation ($z$-)axis lies parallel to the dimer linkage due to the molecular $D_{2d}$ point group symmetry) and this SOC has a significant amplitude of $0.13\mathrm{\,cm^{-1}}$ [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. For comparsion, the nearest-neighbour analogue has a perfect spin selectivity of $\Delta M_{S}=0$ with a larger amplitude of $1.37\mathrm{\,cm^{-1}}$. Note that these SOC matrix elements represent upper bounds to the true molecular values because we have assumed one full electron in each AO.
@@
That ISC in $\pi$-diradicals can actually occur with both spin selectivities is contrary to most theoretical analyses [80, 81, 82] (including ours [52]) and opens up the possibility of transferring the full electronic structure of NV centres into molecules. Our next step is to perfect the next-nearest-neighbour spin selectivity via symmetry considerations. We note that by connecting two benzylic radicals at their para positions, the resulting diradical has at least a $D_{2}$ point group symmetry with irreducible representations (irreps) organised by their characters under $C_{2}\left(y\right)$ and $C_{2}\left(z\right)$ rotations. The irrep of any electronic state is in turn the direct product of its orbital and spin irreps. Focusing on the orbital part, if we continue to align the $z$-axis along the dimer linkage [Fig. Investigating the role of beyond-nearest-neighbour SOCsb,e], then a $C_{2}\left(y\right)$ rotation becomes equivalent to swapping the two monomers. Therefore, for two weakly-interacting $\pi$-systems of significant torsion, the low-lying excitations are expected to appear in near-degenerate pairs of $\left\{A,B_{1}\right\}$ and $\left\{B_{2},B_{3}\right\}$, each comprising a state of opposite parity under $C_{2}\left(y\right)$, i.e. equal probability of exciting either monomer. These two pairs can be distinguished by their transformations under $C_{2}\left(z\right)$ and, for that, it is easier to think in the basis of monomeric $\pi$ MOs. We first note that $2p$ AOs have $\pi$ symmetry, that is, their phases change under $C_{2}$ rotation about the AO centre. Further, for the two $2p$ AOs at the dimer linkage, that $\pi$-symmetry-defining $C_{2}$ rotation coincides with the diradical’s $C_{2}\left(z\right)$ symmetry element. Hence, a monomeric MO with a non-zero amplitude at the dimer linkage must be antisymmetric under $C_{2}\left(z\right)$ rotation. By contrast, for a monomeric MO to be symmetric under $C_{2}\left(z\right)$ rotation, it must possess a node at the dimer linkage. For instance, the Hückel method predicts the highest occupied, singly occupied and lowest unoccupied MOs (abbreviated as HOMO, SOMO, and LUMO respectively) of the benzylic radical to have $+1$, $-1$, and $+1$ characters under $C_{2}\left(z\right)$ rotation [Fig. Investigating the role of beyond-nearest-neighbour SOCsc]. Finally, we recall that the symmetry of the orbital wavefunction is the direct product of the symmetries of all singly occupied orbitals. Therefore, the ground states of benzyl-p-benzyl, having two unpaired electrons in the SOMOs, transform as either $A$ or $B_{1}$. Similarly, the lowest charge transfer (CT) singlets with all doubly-occupied orbitals (also known as charge-recombined or zwitterionic states of the two radicals) also transform as either $A$ or $B_{1}$. As for the lowest local excitations (LEs), they are a linear combination of HOMO-to-SOMO and SOMO-to-LUMO excitations (an outcome of alternancy symmetry; see Ref. [52]) and hence transform as either $B_{2}$ or $B_{3}$. Including the characters under $C_{2}\left(y\right)$ rotation yields the approximate energy level diagram illustrated in Fig. Investigating the role of beyond-nearest-neighbour SOCsd (for a review of the electronic structure of benzylic diradicals, see our earlier work [52]).
@@
Returning to our goal of attaining perfect ISC spin selectivities, we note that the singlet, triplet $M_{S}=0$, and triplet $M_{S}=\pm 1$ spin sublevels transform as $A$, $B_{1}$, and $\left\{B_{2},B_{3}\right\}$ respectively [83, 84]. For an ISC to be symmetry-allowed, the SOC matrix elements must contain the totally symmetric irrep. Hence, because the SOC Hamiltonian is totally symmetric, ISC processes in benzyl-p-benzyl are spin-selective for $\Delta M_{S}=\pm 1$ when moving between triplet LEs and singlet CTs, and $\Delta M_{S}=0$ when moving between singlet CTs and triplet GSs. Actually, we already knew this from the previous section: The first ISC occurs between SOMOs and LUMOs (or HOMOs) of opposite monomers and only the latter has a node at the dimer linkage [Fig. Investigating the role of beyond-nearest-neighbour SOCsc]. Thus, this process occurs predominantly via next-nearest-neighbour SOCs, which we found before to exhibit $\Delta M_{S}=\pm 1$ selectivity [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. Similarly, the second ISC connects between two SOMOs of opposite monomers and is mediated mostly by nearest-neighbour SOCs with $\Delta M_{S}=0$ selectivity [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. Yet, the key improvement achieved by this section is the perfection of ISC spin selectivities to $100\%$ by making the opposite spin channel symmetry-forbidden. From a microscopic perspective, we have destructively interfered any deleterious ISC channels using other higher-nearest-neighbour SOCs, treated more generally by the group-theoretical approach. Importantly, this electronic structure produces the ODMR mechanism shown in Fig. Introductionc, where populations in the triplet $M_{S}=\pm 1$ LEs are moved into the triplet $M_{S}=0$ GS via two sequential ISCs of opposite spin selectivities. In other words, such molecules exhibit the ODMR mechanism of NV centres.
As an aside, while it is tempting to consider only symmetry arguments and ignore the discussions made in the previous section, note that symmetry-allowed transitions need not occur with high amplitudes. Therefore, one still needs to check if the allowed ISCs are facilitated by either nearest- or next-nearest-neighbour SOCs, which have the largest amplitudes. For example, the meta-connected diradicals explored in our previous work [52] have negligible SOCs between singlet CTs and triplet GSs [Fig. Introductionb] because they are dominated by third-nearest-neighbour effects with a maximum SOC amplitude of $0.02\mathrm{\,cm^{-1}}$ [Fig. Investigating the role of beyond-nearest-neighbour SOCsb]. In other words, our plan for enhanced ODMR contrasts requires para-connectivity between the two benzylic radicals, consistent with our group-theoretical analysis (meta-connectivity does not yield $D_{2}$ point group symmetry). We note in passing that vibronic perturbations to the SOC operator, which are not considered by this study, can modify the spin selectivity as well [13, 85]. This will be explored in our upcoming studies.
@@
As speculated, the lowest LE transitions are largely localised on the central (tetra-chlorinated) benzyl-p-benzyl fragment [Fig. Our proposalc]. Hence, for computational efficiency, the SOC matrix elements were computed for this central fragment instead, capping any open ends with hydrogen. Indeed, the MCSCF/CI results suggest appreciable SOC matrix elements along the proposed ODMR pathway [Fig. Our proposalb]. While, strictly speaking, these calculations present only an upper bound to the true SOC matrix elements of the full PT2TM-p-PT2TM diradical, their values should not decrease by much upon weak $\pi$-conjugation to the additional trichlorophenyl substituents. Most importantly, these calculations lend credence to the theoretical proposal presented by this work.
@@
The alternative approach would be to somehow make elements that differ between ${\bf g}_{1}$ and ${\bf g}_{2}$ disappear under experimental conditions. This can be achieved in our diradicals by aligning the magnetic field ${\bf B}$ parallelly to the dimer linkage. (We acknowledge that aligning the molecules, say, in a liquid crystal, may be an experimental challenge, although this has been previously demonstrated [94, 95].) To see that, we first note that elements of the $g$-tensor have the following approximate expression, obtained from second-order perturbation theory in the SOC: [91, 92]
@@
The singlet-triplet states may now block-diagonalise $\hat{H}_{\text{spin}}$ with no singlet-triplet mixings present. Furthermore, due to the approximate $D_{2d}$ point group symmetry of our system, the zero-field splitting caused by ${\bf D}$ also occurs along the $z$-axis. This may be further rationalised by the cylindrical spin density distribution of the diradical [92]. Therefore, the singlet-triplet basis fully-diagonalises $\hat{H}_{\text{spin}}$ in Eq. (23) when the magnetic field is applied along the dimer linkage.
@@
Here, we continue to assume a spin quantisation axis parallel to the dimer linkage. The kinetic processes connecting these eigenstates and their associated rate constants are presented in Fig. Predicted ODMR contrasts; there, the symbols adopt their usual meanings. As far as possible, model parameters were selected within experimental reason. Listed below are our justifications for these parameters and the reader may safely skip it at the first pass.
@@
FIG. 4: : Schematic diagram depicting the excited-state transitions being considered by our microkinetic model.
FIG. 5: : Steady-state solutions to the (a) optically-induced spin polarisation, (b) ODMR contrast, (c) triplet $M_{S}=0$ GS population, and (d) triplet $M_{S}=0$ LE population (which simulates the emission intensity). Plots are made across varying optical pump rates $k_{\text{abs}}$ and spin decoherence rates $k_{\text{decoh}}$ within experimental ranges. Other parameters are defined in Fig. Predicted ODMR contrasts, chosen to be as realistic as possible.